# Integral canonical models of exceptional Shimura varieties

### v1

@@ 2.7 Relation to Hodge theory over $\mathbb{C}$
    If $W_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system ${}_{et}\mathbb{W}_{\ell}$ on $P$, then $V_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system $\mathbb{V}_{\ell}$ on $Y$.
3 Weak stability of Fontaine–Laffaille modules on nodal curves
In this section we prove some semistability properties of Fontaine–Laffaille modules over nodal curves (see Section 3.3). These results will be critical in the main step in the proof of the extension of morphisms from smooth sources in codimenison $\geq 2$ in Section 4.
3.1 The inverse Cartier transform
We begin by recalling the nonabelian Hodge correspondence in characteristic $p$ of Ogus–Vologodsky [OV] via exponential twisting [LSZtwist].
Let $k$ be a perfect field of characterstic $p>2$. We denote by $W_{2}=W_{2}(k)$ the length 2 Witt vectors. Throughout, by a lift of a scheme (usually from $k$ to $W_{2}$) we mean a flat lift. For a variety $Y/k$, we denote by $F_{Y}:Y\to Y$ the absolute Frobenius, dropping the subscript if there’s no chance of confusion. We have a natural lift of Frobenius $F_{W}:W_{2}\to W_{2}$, and in the sequel by a lift of Frobenius to a lift of $Y$ we always mean a lift commuting with $F_{W}$.
3.1.1 Flat sheaves and Higgs sheaves on varieties
We first recall the definitions of Higgs sheaves and flat sheaves on smooth $k$-varieties. As $p\neq 2$ we may unambiguously speak of $\Omega_{X}^{2}$. Throughout we denote by $T_{X}:={\mathcal{H}\operatorname{om}}_{\operatorname{\mathcal{O}}_{X}}(\Omega_{X},\operatorname{\mathcal{O}}_{X})$.
By a flat sheaf on $X$ we mean a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ with a $k$-linear map $\nabla:V\to V\otimes\Omega_{X}$ which satisfies the Leibniz rule and for which the $\operatorname{\mathcal{O}}_{X}$-linear map $\nabla\wedge\nabla:V\to V\otimes\Omega_{X}^{2}$ vanishes. By a Higgs sheaf on $X$ we mean a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ with an $\operatorname{\mathcal{O}}_{Y}$-linear map $\theta:V\to V\otimes\Omega_{X}$ such that $\theta\wedge\theta:V\to V\otimes\Omega^{2}_{X}$ vanishes. Observe that since differentials pull back, both flat sheaves and Higgs sheaves pull back under arbitrary morphisms $f:X\to Y$. We denote by $MIC_{m}(X)$ (resp. $HIG_{m}(X)$) the category of flat sheaves $(V,\nabla)$ (resp. Higgs sheaves $(V,\theta)$) with $p$-curvature $\psi(\nabla):=\nabla^{p}-\nabla_{(-)^{p}}$ (resp. Higgs field $\theta$) nilpotent of exponent $\leq m$, meaning $\psi(\nabla)^{m}=0$ (resp. $\theta^{m}=0$) with the natural notion of morphisms. We will always assume $m\leq p$.
Recall that for any locally free $\mathcal{O}_{X}$-module $V$ on $X$, the pull-back along absolute Frobenius $F^{*}V$ admits a canonical flat connection $\nabla_{can}$ whose flat sections are the pull-back sections $F^{-1}V$. Explicitly, if we write $F^{*}V=\operatorname{\mathcal{O}}_{X}\otimes_{\operatorname{\mathcal{O}}_{X}}V$ where the tensor is taken with respect to the $p$-power algebra structure $\operatorname{\mathcal{O}}_{X}\to\operatorname{\mathcal{O}}_{X}$, then $\nabla_{can}=d\otimes\operatorname{id}$.
3.1.2 Nonabelian Hodge correspondence à la Lan–Sheng–Zuo
Let $Y/k$ be smooth and $\tilde{Y}/W_{2}$ a flat lift. Let $Y_{i}$ be a Zariski cover of $Y$ by affines with overlaps $Y_{ij}:=Y_{i}\cap Y_{j}$ and $\tilde{F}_{i}:\tilde{Y}_{i}\to\tilde{Y}_{i}$ lifts of absolute Frobenius. By lifting to $\operatorname{\mathcal{O}}_{\tilde{Y}}$ and applying $p^{-1}(\tilde{F}_{i}-\tilde{F}_{j})$ (resp. $p^{-1}d\tilde{F}_{i}$) we obtain $F$-derivations of $\operatorname{\mathcal{O}}_{Y_{ij}}$ (resp. $\operatorname{\mathcal{O}}_{Y_{i}}$) valued in $\operatorname{\mathcal{O}}_{Y_{ij}}$ (resp. $\Omega_{Y_{i}}$), which we may think of as $\operatorname{\mathcal{O}}_{Y}$-linear morphisms
$s_{ij}$ $:=p^{-1}(\tilde{F}_{i}-\tilde{F}_{j}):F^{*}\Omega_{Y_{ij}}\to\mathcal{O}_{Y_{ij}}$
$\zeta_{i}$ $:=p^{-1}d\tilde{F}_{i}:F^{*}\Omega_{Y_{i}}\to\Omega_{Y_{i}}$
which satisfy $ds_{ij}=\zeta_{i|j}-\zeta_{j|i}$. For a Higgs sheaf $(V,\theta)\in HIG_{m}(Y)$, it is shown in [LSZtwist] that the inverse Cartier transform $C_{\tilde{Y}}^{-1}(V,\theta)$ is the flat sheaf obtained by locally forming $(F^{*}V|_{Y_{i}},\nabla_{can}+\zeta_{i}(F^{*}\theta))$ on $Y_{i}$ and gluing via
$\exp(s_{ij}(F^{*}\theta)):(F^{*}V|_{Y_{ij}},\nabla_{can}+\zeta_{i}(F^{*}\theta))\to(F^{*}V|_{Y_{ij}},\nabla_{can}+\zeta_{j}(F^{*}\theta))$
up to negating the sign of the connection. Observe that for any Higgs subsheaf $(V^{\prime},\theta^{\prime})\subset(V,\theta)$ we have a natural inclusion of flat sheaves $C_{\tilde{Y}}^{-1}(V^{\prime},\theta^{\prime})\subset C^{-1}_{\tilde{Y}}(V,\theta)$. In particular, if $(G,0)$ is a Higgs subsheaf of $(V,\theta)$, then $(F^{*}G,\nabla_{can})$ is a flat subsheaf of $C^{-1}_{\tilde{Y}}(V,\theta)$.
3.2 An inverse Cartier transform on nodal curves
If $X$ is not smooth, the construction from the previous section fails because a given lift $\tilde{X}$ may not admit local lifts of Frobenius. If $X$ is normal, one may get some mileage out of restricting to the smooth locus as in [langer], and there is a logarithmic version due to Schepler [schepler]. However, the case of interest for us will be that of nodal curves whose nodes smooth out. In this section we construct two types of inverse Cartier transforms: the first for Higgs sheaves pulled-back from a smooth variety $Y$, and the second for Higgs subsheaves of such pull-backs to nodal curves after twisting sufficiently by the ideal of the nodes.
3.2.1 The inverse Cartier transform of pullback Higgs sheaves
Let $\tilde{Y}/W_{2}$ be a smooth $W_{2}$-scheme, $X/k$ a smooth variety, and $f:X\to Y$ a morphism. Let $(V_{Y},\theta_{Y})\in HIG_{m}(Y)$ be a Higgs sheaf on $Y$. Denote $V_{X}=f^{*}V_{Y}$ and by $\rho_{Y|X}:f^{*}\Omega_{Y}\to\Omega_{X}$ the natural map. We have two sorts of Higgs fields on $V_{X}$, each satisfying the obvious integrability condition: $\theta_{Y|X}:V_{X}\to V_{X}\otimes f^{*}\Omega_{Y}$ and $\theta_{X}:V_{X}\to V_{X}\otimes\Omega_{X}$, where $\theta_{X}=(\operatorname{id}\otimes\rho_{Y|X})\circ\theta_{Y|X}$.
@@ 3.2.1 The inverse Cartier transform of pullback Higgs sheaves
The pullback as a flat sheaf $f^{*}C_{\tilde{Y}}^{-1}(V_{Y},\theta_{Y})$ is obtained by locally forming $(F^{*}V_{X},\nabla_{can}+\rho_{Y|X}\circ(f^{*}\zeta_{i})(F^{*}\theta_{Y|X}))$ and gluing via $\exp((f^{*}s_{ij})(F^{*}\theta_{Y|X}))$.
@@
Frobenius commutes with $f$. ∎
In fact the pullback twisting data may be used to define an inverse Cartier transform of a more general kind of object which we call a pullback Higgs sheaf on $X$: precisely, this is a pair $(V,\theta)$ consisting of a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ on $X$ and an $\operatorname{\mathcal{O}}_{X}$-linear map $\theta:V\to V\otimes f^{*}\Omega_{Y}$ satisfying the integrability condition. Indeed, by the same computations as in [LSZtwist], we define a flat sheaf $C^{-1}_{\tilde{Y}|X}(V,\theta)$ by locally forming $(F^{*}V,\nabla_{can}+\rho_{Y\mid X}\circ(f^{*}\zeta_{i})(F^{*}\theta))$ and gluing via $\exp((f^{*}s_{ij})(F^{*}\theta))$. If we define morphisms of pullback Higgs sheaves in the natural way, and $HIG_{m}(Y|X)$ the category of pullback Higgs sheaves of exponent $\leq m$, we obtain the following result, whose details we leave to the reader:
@@
The above definition yields a functor $C^{-1}_{\tilde{Y}|X}:HIG_{p}(Y|X)\to MIC_{p}(X)$ and there is a natural isomorphism $C^{-1}_{\tilde{Y}|X}(V_{X},\theta_{Y|X})\cong f^{*}C^{-1}_{\tilde{Y}}(V_{Y},\theta_{Y})$.
Corollary 3.3.
For $(V_{Y},\theta_{Y})\in HIG_{m}(Y)$ as above, if $(G,0)\to(V_{X},\theta_{Y|X})$ is a morphism of restricted Higgs sheaves, then
$(F^{*}G,\nabla_{can})\cong C^{-1}_{\tilde{Y}|X}(G,0)\to f^{*}C_{\tilde{Y}}^{-1}(V_{Y},\theta_{Y})$
is a morphism of flat sheaves.
3.2.2 The case of nodal curves
Now take $\tilde{X}/W_{2}$ to be a flat nodal curve and $\tilde{\iota}:\tilde{X}\to\tilde{Y}$ a locally closed embedding over $W_{2}$, with base-change $\iota:X\to Y$ to $k$. The case of an embedding will be enough to deduce the case of a general $W_{2}$-liftable morphism (see the proof of Proposition 3.7). Let $\nu:X^{\prime}\to X$ be the normalization of the fiber over $k$, $\iota^{\prime}=\nu\circ\iota$, and let $D$ be the divisor of points of $X^{\prime}$ mapping to nodes of $X$, which we call prenodes. Let $Y_{i}$ be an affine open covering of $Y$. Set $X_{i}=X\cap Y_{i}$ and $X_{ij}=X_{i}\cap X_{j}$; likewise let $X^{\prime}_{i}=\nu^{-1}(X_{i})$ and $X^{\prime}_{ij}=\nu^{-1}(X_{ij})$. We can and will assume that no overlap $X_{ij}$ contains a node.
If $I$ is the ideal of $X$ in $Y$, $N_{X/Y}^{\vee}=I/I^{2}$ is the conormal bundle of $X$, which is locally free. We have a natural exact sequence
[…]
Let $\tilde{I}$ be the ideal of $\tilde{X}$ in $\tilde{Y}$. By lifting to $\tilde{I}$ we have morphisms
$t_{i}:=p^{-1}\tilde{\iota}^{*}\tilde{F}_{i}:F^{*}N_{X_{i}/Y_{i}}^{\vee}\to\mathcal{O}_{X_{i}}.$
since the map $p^{-1}\tilde{F}_{i}:\tilde{I}\to F_{*}\mathcal{O}_{X_{i}}$ kills $pI+\tilde{I}^{2}$. Note that $\rho_{Y|X}\circ\iota^{*}\zeta_{i}\circ F^{*}\delta=dt_{i}$ on $X_{i}$ and $\iota^{*}s_{ij}\circ F^{*}\delta=t_{i|j}-t_{j|i}$ on the overlap $X_{ij}$.
The pullback of (3.2.1) to $X^{\prime}$ fits into a commutative diagram with exact rows and columns
00{0}00{0}00{0}${Q}$${\nu^{*}\Omega_{X}}$${\Omega_{X^{\prime}}}$00{0}00{0}${N_{X^{\prime}/Y}^{\vee}}$${\iota^{\prime*}\Omega_{Y}}$${\Omega_{X^{\prime}}}$00{0}${\nu^{*}N_{X/Y}^{\vee}}$${\nu^{*}N_{X/Y}^{\vee}}$00{0}00{0}$\scriptstyle{\rho_{Y|X^{\prime}}}$$\scriptstyle{\nu^{*}\rho_{Y|X}}$$\scriptstyle{\nu^{*}\delta}$
The locally free sheaf $N_{X^{\prime}/Y}^{\vee}$ is the conormal bundle of the individual branches of $X^{\prime}$ in $Y$. Note that for any locally free sheaf $V$ on a curve and any subspace $U\subset k(x)\otimes V$ of the fiber at $x$, we may define the subsheaf $F$ of $V(x)$ of sections with simple poles in $U$ at $x$ by taking $F(-x)$ to be the kernel of the composition $V\to k(x)\otimes V\to(k(x)\otimes V)/U$. For any prenode $x_{1}$, there is a canonical subspace of the fiber $k(x_{1})\otimes N_{X^{\prime}/Y}$ given by the tangent direction of the curve at the other prenode $x_{2}$ which is attached to $x_{1}$ in $X$. We define $M_{X^{\prime}/Y}\subset N_{X^{\prime}/Y}(D)$ to be the subsheaf of sections of $N_{X^{\prime}/Y}$ with simple poles at each prenode $x_{1}$ in this subspace.

### v2

@@ 2.7 Relation to Hodge theory over $\mathbb{C}$
    If $W_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system ${}_{et}\mathbb{W}_{\ell}$ on $P$, then $V_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system ${}_{et}\mathbb{V}_{\ell}$ on $Y$.
3 Extension in codimension 2
In this section we prove our main extension result in codimension 2.
Theorem 3.1.
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{S}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme equipped with a Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}_{[0,w]}(\hat{\mathcal{S}})$ on the $p$-adic completion $\hat{\mathcal{S}}$ with underlying filtered flat bundle $(M,\nabla,F^{\bullet}M)$ where $20\leq w\leq p-$ 2. Let $\mathcal{X}/\mathcal{O}$ be a smooth finite type $\mathcal{O}$-scheme with $\dim\mathcal{X}=2$, $x\in\mathcal{X}$ a point in the special fiber, and $f:\mathcal{X}\setminus x\to\mathcal{S}$ a morphism. Suppose we have a commutative diagram
${\mathcal{X}^{\prime}}$${\mathcal{S}}$${\mathcal{X}\setminus x}$${\mathcal{X}}$$\scriptstyle{\pi}$$\scriptstyle{g}$
where $\pi:\mathcal{X}^{\prime}\to\mathcal{X}$ is an iterated blow-up at closed points lying above $x$, and $g:\mathcal{X}^{\prime}\to\mathcal{S}$ is a morphism extending $f$ via the natural lift $\mathcal{X}\setminus x\to\mathcal{X}^{\prime}$. Then the Griffiths bundle of $\hat{g}^{*}(M,F^{\bullet}M)$ is trivial on $\pi^{-1}(x)$.
In particular, if the Griffiths bundle of $(M,F^{\bullet}M)$ is ample on $\hat{\mathcal{S}}$, then the morphism $f$ extends over $x$.
We recall the notion of Fontaine–Laffaille modules and the category $\mathrm{MF}_{[0,w]}(\hat{\mathcal{S}})$ in the next section.
3.1 Fontaine-Laffaille modules
3.1.1 Reminder on crystals and flat bundles
Let $F/\mathbb{Q}_{p}$ be an unramified extension and let $\mathcal{O}$ be the ring of integers of $F$. We remind the reader that for a smooth formal scheme $\mathcal{X}/\operatorname{Spf}\mathcal{O}$ there is a natural equivalence of categories between crystals in vector bundles on the special fiber $\mathcal{X}_{0}$ and flat vector bundles on $\mathcal{X}$ (see for example [stacks-project, Tag 07JH]). In particular, there is a well-defined “Frobenius pull-back” of a flat vector bundle $(V,\nabla)$ on $\mathcal{X}$ which we denote $F^{*}(V,\nabla)$. Concretely, for a lift of Frobenius $\phi:\mathcal{X}\to\mathcal{X}$, we form $(\phi^{*}V,\phi^{*}\nabla)$, and for any other lift of Frobenius $\phi^{\prime}:\mathcal{X}\to\mathcal{X}$ there is a canonical isomorphism of flat vector bundles $(\phi^{\prime*}V,\phi^{\prime*}\nabla)\to(\phi^{*}V,\phi^{*}\nabla)$ given by the formula:
$1\otimes v\mapsto\sum_{I}\frac{(\phi^{\prime}(t)-\phi(t))^{I}}{I!}\otimes\nabla(\partial)^{I}v.$ (3.1.1)
Here we take $t_{1},\dots,t_{n}$ to be local co-ordinates for $\mathcal{X}$ with a dual basis of derivations $\partial_{i}$ and for a multi-index $I=(i_{1},\dots,i_{n})$ we set $\nabla(\partial)^{I}=\prod_{j}\nabla(\partial_{i})^{i_{j}}$.
The above isomorphisms satisfy an appropriate cocycle condition, hence local pullbacks (where a lift of Frobenius exists) canonically glue. Moreover, for any morphism $f:\mathcal{Y}\to\mathcal{X}$ of smooth formal schemes there is a natural identification $f^{*}F^{*}(V,\nabla)\cong F^{*}f^{*}(V,\nabla)$ by factoring $f$ as the graph followed by the projection $\mathcal{X}\to\mathcal{X}\times\mathcal{Y}\to\mathcal{Y}$ and using the existence of campatible lifts of Frobenius for the graph and the projection separately.
3.1.2 Unramified case
Let $F/\mathbb{Q}_{p}$ be an unramified extension and let $\mathcal{O}$ be the ring of integers of $F$. Let $\mathcal{X}$ denote a smooth formal scheme over $\mathcal{O}$, and $\mathcal{X}^{rig}$ its rigid-analytic fiber. In [faltings] Faltings defines a category of Fontaine-Laffaille modules as follows:
Definition 3.2.
Assume first that $\mathcal{X}$ admits a lift of Frobenius $\phi:\mathcal{X}\rightarrow\mathcal{X}$. Let $20\leq w\leq p-$ 2. Then an object of $\operatorname{MF}_{[0,w]}(\mathcal{X})$ consists of a vector bundle $M$ equipped with a decreasing filtration $F^{\bullet}M$ with $F^{0}M=M,F^{a+1}M=0$, a Griffiths-transverse flat connection $\nabla_{M}$, and a morphism $\phi_{M}:\phi^{*}M\rightarrow M$ satisfying:
  1. 1.
    $\phi_{M}$ is horizontal for $\nabla_{M}$
  2. 2.
    The restriction of $\phi_{M}$ to $\phi_{M}^{*}F^{q}M$ is divisible by $p^{q}$, for $q\in[0,w]$.
  3. 3.
    $M=\sum_{i=0}^{w}p^{-i}\phi_{M}(\phi_{M}^{*}F^{i}M)$.
Morphisms in $\operatorname{MF}_{[0,w]}(\mathcal{X})$ are morphisms of the underlying modules which are compatible the additional structures.
Faltings shows that this category is independent of the choice of $\phi$ up to a natural isomorphism, using the canonical identifications of eq. 3.1.1. Indeed, for a different choice of Frobenius lift $\phi^{\prime}:\mathcal{X}\rightarrow\mathcal{X}$, we obtain a new lift $\phi_{M}^{\prime}$ to $M$ by setting:
$\phi^{\prime}_{M}(1\otimes m):=\sum_{I}\phi(1\otimes\nabla_{M}(\partial^{I})(m))\frac{(\phi^{\prime}(t)-\phi(t))^{I}}{I!}.$ (3.1.2)
Finally, for a general $\mathcal{X}$, one simply covers by affines where a lift of Frobenius exists as above and glues in the natural way. Since Frobenius lifts lift uniquely under etale maps, it follows that $\operatorname{MF}_{[0,w]}(\mathcal{X})$ satisfies étale descent.
Whenever we write $\operatorname{MF}_{[0,w]}$ we assume that $20\leq w\leq p-$ 2. The following lemma must be well-known to the experts but we cannot find it in the literature in this form:
@@ 3.1.2 Unramified case
Let $g:\mathcal{X}\rightarrow\mathcal{Y}$ be a map of smooth formal $\mathcal{O}$-schemes. There is a natural base-change morphism $\operatorname{MF}(g):\operatorname{MF}_{[0,w]}(\mathcal{Y})\rightarrow\operatorname{MF}_{[0,w]}(\mathcal{X})$, compatible with compositions, so that $\operatorname{MF}(g_{1}\circ g_{2})=\operatorname{MF}(g_{1})\circ\operatorname{MF}(g_{2}).$
@@
The statement is clearly Zariski-local on both $\mathcal{X}$ and $\mathcal{Y}$, which we may therefore assume to be affine. We may simply pull back the vector bundle, filtration, and connection. The only tricky part lies in the semi-linear Frobenius morphism $\phi$. As explained in [faltings], the data of $(M,\nabla_{M},\phi_{M})$ is equivalent to an $F$-crystal on $\mathcal{X}_{0}$. Since $F$-crystals pull back naturally, the claim is proven.
3.1.3 Ramified case
Let $F^{\prime}/\mathbb{Q}_{p}$ be a finite extension, which might be ramified. Let $F\subset F^{\prime}$ be the maximal unramified subfield, and let $\mathcal{O}^{\prime},\mathcal{O}$ be the rings of integers of $F^{\prime},F$. Let $\pi\in\mathcal{O}$ denote a uniformizer, with minimal polynomial $f(t)\in\mathcal{O}[t]$. Let $R=\mathcal{O}[[t]][\frac{f^{n}}{n!}]$, and let $R_{\mathcal{O}^{\prime}}$ be the PD-completion of $R$. Then $R_{\mathcal{O}^{\prime}}$ naturally acquires a decreasing filtration with $F^{0}R_{\mathcal{O}^{\prime}}=R_{\mathcal{O}^{\prime}}$ and $F^{q}R_{\mathcal{O}^{\prime}}$ being generated by the divided powers $\frac{f^{n}}{n!},n\geq q$.
In [faltingsramified, Def 2] Faltings defines a category $\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$ for $a\leq p-2$ whose definition we recall here:
Definition 3.4.
An object $\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$ is a $R_{\mathcal{O}^{\prime}}$-module $M$ with a decreasing filtration $F^{\bullet}M$, a Griffiths-transverse flat connection $\nabla_{M}$ which nilpotent modulo $p$, and a (semi-linear) Frobenius $\phi_{M}$ satisfying the following:
  1. 1.
    $M$ is filtered-free as an $R_{\mathcal{O}^{\prime}}$-module (meaning a direct sum of twists of $R_{\mathcal{O}^{\prime}}$), with a basis $m_{i}$ in degrees $q_{i}$ in $[0,w]$.
  2. 2.
    $\phi_{M}$ is horizontal for $\nabla_{M}$
  3. 3.
    The restriction of $\phi$ to $F^{q}M$ is divisible by $p^{q}$, for $q\in[0,w]$.
  4. 4.
    The elements $\phi(m_{i})/p^{q_{i}}$ form another basis for $M$.
Morphisms in $\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$ are morphisms of the underlying modules which are compatible the additional structures.
Once again, as Faltings observes, the data of $(M,\nabla_{M},\phi)$ gives an $F$-crystal on $\mathcal{O}^{\prime}\otimes\mathbb{F}_{p}$.
@@ 3.1.3 Ramified case
Let $\mathcal{S}$ be a smooth formal scheme over $\operatorname{Spf}\mathcal{O}$, and let $\iota:\operatorname{Spf}\mathcal{O}^{\prime}\rightarrow\mathcal{S}$ be a morphism of formal $\mathcal{O}$-schemes. There is a natural base-change morphism $\operatorname{MF}(\iota):\operatorname{MF}_{[0,w]}(\mathcal{S})\rightarrow\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$.
Proof.
Let $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{S})$. We may define the flat bundle $(M_{1},\nabla_{M_{1}})$ underlying $\operatorname{MF}(\iota)(\mathbf{M})$ and the Frobenius $\phi_{M_{1}}$ by simply pulling back the $F$-crystal underlying $M$ to $\operatorname{Spec}\mathcal{O}^{\prime}$. It remains to define the filtration. Without loss of generality, we may assume $\mathcal{S}$ is affine, so $\iota:A\rightarrow\mathcal{O}^{\prime}$ for $A$ the p-adic completion of a smooth $\mathcal{O}$-algebra. By smoothness, we may pick a lift $\iota_{1}:A\rightarrow R_{\mathcal{O}^{\prime}}$, and note $M_{1}\cong R_{\mathcal{O}^{\prime}}\otimes_{A}M$. We now get a natural filtration on $M_{1}$ by taking the product filtration. We must show that this is independent of the choice of $\iota_{1}$, so let $\iota_{2}$ be a different lift. Then the crystal structure gives an isomorphism of flat bundles $M_{1}\rightarrow M_{2}$ given as follows. As before, let $t_{1},\dots,t_{n}$ be local co-ordinates for $X$ with a dual basis of derivations $\partial_{i}$. For a multi-index $I=(i_{1},\dots,i_{n})$ set $\nabla_{M}(\partial)^{I}=\prod_{j}\nabla_{M}(\partial_{i})^{i_{j}}$.
Then our isomorphism $M_{1}\rightarrow M_{2}$ is given by:
$1\otimes_{\iota_{1}}m\mapsto\sum_{I}\frac{(\iota_{1}(t)-\iota_{2}(t))^{I}}{I!}\otimes_{\iota_{2}}\nabla_{M}(\partial^{I})(m).$
Since $\frac{(\iota_{1}(t)-\iota_{2}(t))^{I}}{I!}\in F^{|I|}R_{\mathcal{O}^{\prime}}$ and the connection is Griffiths transverse, it follows that this isomorphism respects filtrations. This completes the proof
We point out that if $C_{\mathbf{M}}$ denotes the crystal corresponding to $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$, then $C_{\mathbf{M}}(\operatorname{Spf}\mathcal{O}^{\prime})\cong\mathcal{O}^{\prime}\otimes_{R_{\mathcal{O}^{\prime}}}M$.
3.1.4 Local systems
Following Faltings, we say that a small $\mathcal{O}$-algebra is one that admits a finite, etale map from $\mathcal{O}\langle T_{1}^{\pm},\dots,T_{n}^{\pm}\rangle$. Let $\mathcal{S}$ be a smooth formal $\mathcal{O}$-scheme. Using the above notion of Fontaine–Laffaille module, in [faltingsramified, Thm 5] there is defined a fully faithful functor $\mathbb{D}:\operatorname{MF}_{[0,w]}(\mathcal{S})\rightarrow\operatorname{Loc}_{S^{rig}}(\mathbb{Z}_{p})$. We recall the notion here, and to do so we recall some basic rings that are used:
Let $\mathcal{S}=\operatorname{Spf}A$ where $A$ is a small $\mathcal{O}$-algebra. Consider the maximal extension of the fraction field $\operatorname{Frac}(A)$ in which $A[1/p]$ has unramified normalization, and let $\overline{A}$ be the normalization of $A$ in this field. We let $A^{\flat}$ be the perfection of $\overline{A}/p$ and $W(A^{\flat})$ the Witt ring. Then there is a map $W(A^{\flat})\rightarrow\widehat{\overline{A}}$, and we let $B^{+}(A)$ denote the completed divided power hull with respect to the ideal generated by the kernel $I$ of this map and $p$. The group $G_{A}:=\operatorname{Gal}(\overline{A}[1/p]/A[1/p])$ acts continuously on $B^{+}(A)$.
Giving an $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{S})$ gives an $F$-crystal $C_{\mathbf{M}}$ on $\mathcal{S}_{0}$, and hence by pull back an $F$-crystal $\overline{C}_{\mathbf{M}}$ on $\widehat{\overline{A}}\otimes\mathbb{F}_{p}$ with a $\operatorname{Gal}(\overline{A}/A)$ action. As in lemma 3.5, this endows $\overline{C}_{\mathbf{M}}(B^{+}(A))$ with a natural filtration. We define
$\mathbb{D}(\mathbf{M}):=\operatorname{Hom}_{B^{+}(A),{\mathrm{Fil}},\phi}(\overline{C}_{\mathbf{M}}(B^{+}(A)),B^{+}(A))$
where the homomorphism is of $B^{+}(A)$-modules and is required to respect the filtration and Frobenius.
The analogous definition using $B^{+}(\mathcal{O}^{\prime})$ gives a functor $\mathbb{D}:\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})\rightarrow\operatorname{Loc}_{\mathcal{O}^{\prime}}(\mathbb{Z}_{p})$, where $\operatorname{Loc}_{\mathcal{O}^{\prime}}(\mathbb{Z}_{p})$ simply denotes free $\mathbb{Z}_{p}$ representations of $\operatorname{Gal}_{F^{\prime}}$.